# Random divergence-free drifts and the Onsager-Richardson threshold

### 12 Mar 2026 v1

@@ 3. Proof of the Morse-Sard property
On the other hand, since $B_{C_{0}r_{i}}(z_{i})\cap V$ is a closed (d2)(d-2)-dimensional disk, the set $W_{i}$ can be covered by $N_{i}\leq\bar{C}r_{i}^{-\alpha(d-2)}$ balls (of $\mathbb{R}^{d-1}$) with radius $r_{i}^{1+\alpha}$, where $\bar{C}$ depends only on $C_{0}$, $D\phi]_{0,\alpha}$, and $d$. The union over $i$ of all these balls covers therefore $\phi(Z)$ and we can write
$|\phi(Z)|$ $\leq\omega_{d-1}\sum_{i}N_{i}(r_{i}^{1+\alpha})^{d-1}\leq\bar{C}\omega_{d-1}\sum_{i}r_{i}^{(1+\alpha)(d-1)-\alpha(d-2)}$
$=\bar{C}\omega_{d-1}\sum_{i}r_{i}^{d-1+\alpha}\,.$ (3.14)
Observe that $d-1+\alpha\geq d-2\alpha+\delta$ because of nequality (3.11). Moreover, since $\varepsilon\leq 1$, inequality (3.13) implies that $r_{i}\leq 1$ for all $i$’s and in particular $r_{i}^{d-1+\alpha}\leq r_{i}^{d-2\alpha+\delta}$. Thus (3.13) and (3.14) immediately imply
$|\phi(Z)|\leq\bar{C}\omega_{d-1}\varepsilon\,.$
Since $\varepsilon$ can be chosen arbitrarily small, we then conclude that $\phi(Z)$ has zero Lebesgue measure, which is what we had to show. ∎
4. Remarks
In this section we take Theorem 1.2 as base of our discussions. Analogous statements hold for Theorem 1.3 and Theorem 2.2, but we leave the details to the reader.
4.1. Random vector fields
First of all, we observe that assumption (b) in Theorem 1.2 is equivalent to the probability distribution of the ($\mathbb{R}^{2}$-valued) random variable $v(x)$ having bounded density. This is an extremely mild assumption, which is satisfied in a number of standard choices of $\mathbb{P}$. Consider for instance the usual construction of random vector fields as follows
[…]
Assumption (b) amounts therefore to having a uniform lower bound $C(x)\geq C_{0}{\rm Id}$ in the sense of positive definite matrices. If we take classical bases such as Fourier or (several bases of) wavelets, the uniform lower bound is implied by the positivity of the $\sigma_{k}$’s. In fact the positivity of them for a handful of low frequencies is already enough.
Furthermore, a simple inspection of the proof shows that indeed condition (b) is only needed for the special value $v_{0}=0$, while in the general case of Theorem 1.3 the condition (b) is only needed for $M_{0}$’s which have rank at most d2d-2.
4.2. Adding constant drifts
The assumption, in Theorem 1.2, that the average $\langle v\rangle$ of the random vector field $v$ vanishes almost surely can be easily removed. Here we quickly explain how. First of all, the arguments of Lemma 3.1 can be used verbatim to show that, $\mathbb{P}$-almost surely, the set $\{v=0\}$ has dimension at most $2-2\alpha$: the proof only uses the assumptions (a) and (b). Fix now such a $v$. The claim is that it satisfies the conclusions of the theorem because in fact it has the DiPerna-Lions renormalization property: the assumption $\langle v\rangle=0$ is not essential in this step either. First of all Lemma 3.2 can be used verbatim to show that, in every subset $\Omega$ where a function $\psi$ with $\nabla\psi^{\perp}=v$ exists, the critical values of $\psi$ has measure zero. Since this will hold for every $\Omega$ which is simply connected, the results in [2] can then be used to show that, when $\Omega$ is simply connected and $\theta$ and $\beta$ are as in Definition 2.1, then $\beta(\theta)$ will be a distributional solution of the transport equation (2.1) in $\Omega\times[0,\infty)$ with initial data $\beta(\theta_{in})$. But a simple partition of unity allows then to show that $\beta(\theta)$ is in fact a distributional solution on the whole torus.
Generalizing the above to Theorem 2.2 is rather straightforward: we just need to make $\mathbb{P}$ a probability distribution on the space of $C^{1}$ maps $\phi:\mathbb{R}^{d}\to\mathbb{R}^{d-1}$ which are the sums of a $1$-periodic function and a linear map. The differential of each component $\phi_{i}$ is then a periodic $1$-form and induce therefore a closed $1$-form $\omega_{i}$ on $\mathbb{T}^{d}$. In particular the conclusion of Theorem 2.2 will then hold for the divergence-free vector field $w$ (which needs not be mean-free) such that $\star\sum_{i}w_{i}dx_{i}=\omega_{1}\wedge\ldots\wedge\omega_{d-1}$.
Remark 4.1.
The addition of a sufficiently large constant to a $2$-dimensional $C^{1}$ divergence-free autonomous vector field makes it a nowhere zero vector field. It follows then immediately from our discussion that $w$ has the renormalization property: this of course a very simple case of the general theory developed in [2]. Consider now a probability measure $\mathbb{P}$ on divergence-free autonomous vector fields with the following two properties:
  • $v\in C^{0}$ $\mathbb{P}$-almost surely;
  • The average $\langle v\rangle$ of $v$ and its average-free part $v-\langle v\rangle$ are independent;
  • $\mathbb{P}(\{v:|\langle v\rangle|\geq M\})>0$ for every constant $M$.
Then there is a nonzero probability $\mathbb{P}$ that $v$ has no zeros, and thus a nonzero probability that it has the renormalization property. Note that the probability is, potentially, rather small but that no Hölder regularity of $v$ is needed. Note that the fact that in Theorems 1.2, 1.3 and 2.2 we restrict to zero average vector fields means that they have to be zero somewhere.
4.3. Yaglom’s law
A widely accepted theory in the turbulence literature is that there is, in some statistical sense, a relation between the regularity of the advecting vector field $v$ and the passive scalar $\theta_{\varepsilon}$ in (1.5). Roughly speaking a $C^{\alpha}$ regularity for $v$ should correspond to a uniform $C^{\beta}$ regularity of $\theta_{\varepsilon}$ for $\beta$ satisfying $\alpha+2\beta=1$. Some authors in the literature call this “anomalous regularization”. More generally the “law” is known as Yaglom’s Law or the Obukhov-Corrsin relation. Clearly a deterministic statement of the latter cannot hold. However, Theorem 1.2 (and, to a lesser extent, also Theorem 2.2) show that even at a statistical level such a statement cannot hold when $v$ is autonomous and $\alpha>\frac{1}{3}$. In fact, because of the DiPerna-Lions renormalization property, when the conclusion of Theorem 2.2 applies and the initial data $\theta_{in}$ is bounded, the solutions $\theta_{\varepsilon}$ of (1.5) converge (strongly in $C([0,T],L^{2})$) to the corresponding unique solution $\theta$ of (2.1). Moreover $\theta$ is constant along the trajectories of the unique regular Lagrangian flow $\Phi$ of the vector field $v$. Therefore the distribution function of $a\mapsto|\{\theta>a\}|$ will remain exactly the same. This immediately obstructs any regularization mechanism if, for instance, the values taken by $\theta_{in}$ is a discrete set with more than one element: $\theta(\cdot,t)$ will be forced to remain discontinuous on a set of positive $\mathcal{H}^{1}$ measure.
Additionally, Remark 4.1 suggests that a proper framework to study random drifts is likely that of average zero vector fields.
@@ 4.4. From the renormalization property to the absence of dissipation anomalies
First of all the existence and uniqueness of the bounded distributional solution $\theta_{0}$ of (2.1) is the main result of the DiPerna-Lions theory and the reader can consult e.g. [20, Theorem 3.3 & Proposition 3.6]. The strong continuity of $t\mapsto\theta(\cdot,t)$ (up to a suitable choice of representative) is another well-known fact and the reader can consult e.g. [4, Remark 25]. By the maximum principle $\theta_{\varepsilon}$ is uniformly bounded in $L^{\infty}((0,T);L^{\infty}(\mathbb{T}^{d}))$ and so, up to subsequence, it converges weakly-* to some bounded function $\theta$. We immediately see (passing to the limit in the weak formulation (2.2)) that $\theta$ is a bounded distributional solution of equation (2.1) and the aforementioned uniqueness implies that $\theta=\theta_{0}$. Fix a smooth test function $\varphi\in C^{\infty}(\mathbb{T}^{d})$ and define $\Phi_{\varepsilon}(t):=\int_{\mathbb{T}^{d}}\varphi(x)\theta_{\varepsilon}(x,t)dx$. Multiplying equation (1.5) with the function $\varphi(x)$ and integrating in space, we immediately obtain that

### 01 Apr 2026 v2

@@ 2.4. Absence of triods
On the other hand, since $B_{C_{0}r_{i}}(z_{i})\cap V$ is a closed $k$-dimensional disk, the set $W_{i}$ can be covered by $N_{i}\leq\bar{C}r_{i}^{-\alpha k}$ balls (of $\mathbb{R}^{d-1}$) with radius $r_{i}^{1+\alpha}$, where $\bar{C}$ depends only on $C_{0}$, $D\phi]_{0,\alpha}$, and $d$. The union over $i$ of all these balls covers therefore $\phi(Y)$ and, if we fix $\gamma>0$, we can write
$\mathcal{H}^{\gamma}_{\infty}(\phi(Y))$ $\leq C\sum_{i}N_{i}(r_{i}^{1+\alpha})^{\gamma}\leq\bar{C}\sum_{i}r_{i}^{(1+\alpha)\gamma-\alpha k}$ (2.17)
Choose now $\gamma=\frac{\beta+\alpha k+\delta}{1+\alpha}$ so that
$(1+\alpha)\gamma-\alpha k=\beta+\delta$
In particular we have, by inequality (2.16),
$\mathcal{H}^{\gamma}_{\infty}(\phi(Y))\leq\bar{C}\varepsilon\,$,
where the constant $\bar{C}$ depends on $\phi$, $\beta$, and $\delta$ but not on $\varepsilon$. In particular letting $\varepsilon\downarrow 0$ we conclude that
$\mathcal{H}^{\frac{\beta+\delta+k\alpha}{1+\alpha}}(\phi(Y))=0$
On the other hand, given that $\delta>0$ is arbitrary, this proves that the Hausdorff dimension of $\phi(Y)$ is no larger than $\frac{\beta+k\alpha}{1+\alpha}$.
3. Consequences of the DiPerna-Lions theory
As mentioned in the introduction, we collect here a number of consequences of the DiPerna-Lions theory. The focus is to show that a vector field which satisties the DiPerna-Lions renormalization property necessarily violates many of the expected “laws” of turbulent fluids. This seems to be folklore in the literature, but we believe it is useful to give precise statements and include the corresponding arguments.
@@ 3.1. From the renormalization property to the absence of dissipation anomalies
First of all the existence and uniqueness of the bounded distributional solution $\theta_{0}$ of (2.1) is the main result of the DiPerna-Lions theory and the reader can consult e.g. [27, Theorem 3.3 & Proposition 3.6]. The strong continuity of $t\mapsto\theta(\cdot,t)$ (up to a suitable choice of representative) is another well-known fact and the reader can consult e.g. [5, Remark 25]. By the maximum principle $\theta_{\varepsilon}$ is uniformly bounded in $L^{\infty}((0,T);L^{\infty}(\mathbb{T}^{d}))$ and so, up to subsequence, it converges weakly-* to some bounded function $\theta$. We immediately see (passing to the limit in the weak formulation (2.2)) that $\theta$ is a bounded distributional solution of equation (2.1) and the aforementioned uniqueness implies that $\theta=\theta_{0}$. Fix a smooth test function $\varphi\in C^{\infty}(\mathbb{T}^{d})$ and define $\Phi_{\varepsilon}(t):=\int_{\mathbb{T}^{d}}\varphi(x)\theta_{\varepsilon}(x,t)dx$. Multiplying equation (1.5) with the function $\varphi(x)$ and integrating in space, we immediately obtain that